solve the following equations 2x^2-x=21 show work
Answer by mananth(16946)

2x^2-x=21
2x^2-x-21=0
2x^2-7x+6x-21=0
x(2x-7)+3(2x-7)=0
(2x-7)(x+3)=0
x=7/2 OR -3